# TimeBridge: Better Diffusion Prior Design with Bridge Models for Time Series Generation

### v2

@@ Appendix A Diffusion Bridge
We introduce the details of the diffusion bridge mathematically, and how this framework is embraced in recent deep learning architectures.
@@ A.1. Schrödinger Bridge
Standard diffusion models suffer from limitations such as the requirement of a sufficiently long time for the prior distribution to approximate a standard Gaussian distribution and the computational expense associated with sample generation. To address these challenges, diffusion Schrödinger bridge models have been recently introduced (De Bortoli et al., 2021; Chen et al., 2021). The goal of the Schrödinger bridge problem is to find an entropic optimal transport between two probability distributions in terms of Kullback–Leibler divergence on path spaces within a finite horizon (Chen et al., 2021; Léonard, 2012). While standard diffusion models and flow matching models are not guaranteed to provide optimal transport, Schrödinger bridge models aim to find paths that recover entropy-regularized versions of optimal transport (Shi et al., 2024). The formulation of the Schrödinger bridge problem is as follows:
@@
where $\mathcal{P}_{[0,T]}$ refers to the space of path measures on $[0,T]$, $p^{ref}$ denotes the reference path measure, and $p_{0},p_{T}$ represents the marginal distributions of $p$ at each time step $0,T$, respectively. The diffusion bridge algorithm is introduced by utilizing $p^{ref}$ from the forward SDE in Equation (4) (Vargas et al., 2021; Chen et al., 2021; De Bortoli et al., 2021). Setting the reference path as the forward SDE, the problem becomes equivalent to following forward-backward SDEs:
(23) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\\ \quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(24) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\\ \quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are described by following PDEs:
@@
(25) $\begin{split}\frac{\partial\bm{\Psi}}{\partial t}&=-\nabla_{x}\bm{\Psi}^{T}{\mathbf{f}}-\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\bm{\Psi})\\ \frac{\partial\hat{\bm{\Psi}}}{\partial t}&=-\nabla_{x}\cdot(\hat{\bm{\Psi}}{\mathbf{f}})+\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\hat{\bm{\Psi}})\\ &s.t.\quad\bm{\Psi}_{0}\hat{\bm{\Psi}}_{0}=p_{data},\bm{\Psi}_{T}\hat{\bm{\Psi}}_{T}=p_{prior},\\ &\qquad\quad p_{t}=\bm{\Psi}_{t}\hat{\bm{\Psi}}_{t}\ \text{for}\ t\in[0,T].\end{split}$
@@
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940). At the outset, several early studies employed iterative methods to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). A classical solver for Schrödinger bridges is Iterative Proportional Fitting (IPF). Beginning with arbitrary half-potentials $(\bm{\Psi}^{(0)}_{t},\hat{\bm{\Psi}}^{(0)}_{t})$, the algorithm alternates two normalization steps:
[…]
@@
After each full iteration, the product $\bm{\Psi}^{(n+1)}_{t}\hat{\bm{\Psi}}^{(n+1)}_{t}$ matches both marginals, the sequence converges geometrically, requiring one forward–backward integration per iteration. However, these methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
@@ A.2. Denoising Diffusion Bridge Models
Alternatively, Chen et al. (2021) leveraged the theory of forward-backward SDEs to derive an exact log-likelihood expression for the Schrödinger bridge, which accurately generalizes the approach for score generative models. Drawing inspiration from the perspective that the framework of diffusion bridges, employing forward-backward SDEs, encompasses the paradigms of score matching diffusion models (Song et al., 2020) and flow matching optimal transport paths (Lipman et al., 2022). Zhou et al. (2024) demonstrated that reintroducing several design choices from these domains becomes feasible. In particular, the reparameterization techniques outlined in Karras et al. (2022) are utilized.
Table 9. VP and VE instantiations of diffusion bridge settings in DDBM.
@@
For noise scheduling, common options include variance-preserving (VP) and variance-exploding (VE) diffusion (Song et al., 2020). We express the reparameterization for the VP and VE bridge of Zhou et al. (2024) in Table 9. In our study, we choose to utilize the VP bridge, as our experimental findings indicate its superiority of VP over VE.
For score matching, Karras et al. (2022) proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected $L_{2}$ error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we can train $D_{\theta}$ directly instead of training the score function. In general, training a neural network to model $D_{\theta}$ directly is ineffective, due to its vast variance, which heavily depends on the noise level. Thus, Karras et al. (2022) reconstructed $D_{\theta}$ with $\sigma$-dependent skip connection as follows:
(28) $D_{\theta}({\mathbf{x}};\sigma)=c_{skip}(\sigma){\mathbf{x}}+c_{out}(\sigma)F_{\theta}(c_{in}(\sigma){\mathbf{x}};c_{noise}(\theta)),$
and trained $F_{\theta}$ instead of $D_{\theta}$ with the neural network. However, for time series data, estimating the diffusion noise poses greater challenges due to the presence of highly irregular noisy components (Shen and Kwok, 2023; Yuan and Qiao, 2024). Consistent with these studies, our experiments demonstrate that $D$-matching is more effective than $F$-matching. Detailed experimental results for comparison of VP with VE and $D$-matching with $F$-matching are provided in the ablation study in the Appendix C.6.
@@ Appendix B Experimental Settings
B.1. Situations
(a) Unconditional
(b) Linear trend
(c) Polynomial trend
(d) Fixed point
(e) Missing value
B.2. Baseline methods
For the unconditional baseline methods, we compare with TimeVAE (Desai et al., 2021), TimeGAN (Yoon et al., 2019), Cot-GAN (Xu et al., 2020), Diffwave (Kong et al., 2021), DiffTime (Coletta et al., 2024), and Diffusion-TS (Yuan and Qiao, 2024). For Table 2, we note the quality evaluation results in Diffusion-TS. We reimplement Diffusion-TS using their official GitHub in https://github.com/Y-debug-sys/Diffusion-TS, unfortunately, we cannot achieve the performance of Diffusion-TS in their original paper.
Furthermore, for the conditional generations, we reimplement SSSD (Alcaraz and Strodthoff, 2023) using the official GitHub https://github.com/SSSD/sssd. For the trend-guided synthesis, we treat the trend as conditional observations and train the SSSD model to match the total time series data. We add the conditional embedding layers in their original model architectures.
B.3. Evaluation Metrics
FID score is a widely used metric in image generation tasks, capturing the similarity between generated images and real ones (Heusel et al., 2017). To adapt it to the time series domain, Jeha et al. (2022) proposed Context-FID score, which replaces the Inception model used in the original FID with TS2Vec, a pre-trained time series representation learning model (Yue et al., 2022). As Context-FID score aligns with the predictive accuracy of the generative model in downstream tasks, a lower Context-FID score indicates higher synthetic data quality. For computation, first, we sample generated synthetic time series and real-time series separately. Next, both are encoded using a pre-trained TS2Vec model, and the FID score is calculated based on the resulting representations.
Correlational score is a metric used to evaluate the similarity in the covariance of time series features between original data and synthetic data (Liao et al., 2020). First, we estimate the covariance between the $i$-th and $j$-th features of the time series by:
Then compute the correlation between the original data and synthetic data as follows:
B.4. Additional Experimental Details
Datasets
Table 10. Details of datasets.
Dataset Sample size # of features Datatset link
Sines 10000 5 https://github.com/jsyoon0823/TimeGAN
Stocks 3773 6 https://finance.yahoo.com/quote/GOOG
ETTh 17420 7 https://github.com/zhouhaoyi/ETDataset
MuJoCo 10000 14 https://github.com/google-deepmind/dm_control
Energy 19711 28 https://archive.ics.uci.edu/dataset
fMRI 10000 50 https://www.fmrib.ox.ac.uk/datasets
Diffusion Sampler
We further explain the details of our sampler for the unconditional situation in Algorithm 2. For implementing the diffusion sampler, we modify the official code of DDBM (Zhou et al., 2024) in https://github.com/alexzhou907/DDBM as the backbone of our sampler.
Input: Diffusion model $D_{\theta}({\mathbf{x}}_{t})$, score function ${\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i}},t_{i},{\mathbf{x}}_{T},{T})$ defined in Equation 12, diffusion steps $\{0=t_{0}<\cdots<t_{\Gamma}=T\}$, step ratio $s$
Output: Denoised output ${\mathbf{x}}_{0}$
Sample ${\mathbf{x}}_{T}\sim\textit{prior}$ using Equations (19) or (20)
for $i=\Gamma,\ldots,1$ do
       Sample $\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
       $\tilde{t}_{i}\leftarrow t_{i}+s(t_{i-1}-t_{i})$
       $\tilde{{\mathbf{x}}}_{t_{i}}\leftarrow{\mathbf{x}}_{t_{i}}+{\mathbf{d}}_{t_{i}}(\tilde{t}_{i}-t_{i})+g(t_{i})\sqrt{\tilde{t}_{i}-t_{i}}\cdot\bm{{\bm{\varepsilon}}}$
       $\tilde{{\mathbf{d}}}_{t_{i}}\leftarrow-{\mathbf{f}}(\tilde{{\mathbf{x}}}_{\tilde{t}_{i}},\tilde{t}_{i})+g^{2}(\tilde{t}_{i})(\frac{1}{2}{\mathbf{s}}(D_{\theta},\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},{T})-{\mathbf{h}}(\tilde{{\mathbf{x}}}_{t_{i}},\tilde{t}_{i},{\mathbf{x}}_{T},T))$
       ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\tilde{{\mathbf{d}}}_{t_{i}}(t_{i-1}-\tilde{t}_{i})$
       if $i\neq 1$ then
             ${\mathbf{d}}^{\prime}_{t_{i}}\leftarrow-{\mathbf{f}}({\mathbf{x}}_{t_{i-1}},t_{i-1})+g^{2}(t_{i-1})(\frac{1}{2}{\mathbf{s}}(D_{\theta},{\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},{T})-{\mathbf{h}}({\mathbf{x}}_{t_{i-1}},t_{i-1},{\mathbf{x}}_{T},T))$
            ${\mathbf{x}}_{t_{i-1}}\leftarrow\tilde{{\mathbf{x}}}_{t_{i}}+\frac{1}{2}({\mathbf{d}}^{\prime}_{t_{i}}+\tilde{{\mathbf{d}}}_{t_{i}})(t_{i-1}-\tilde{t}_{i})$
Algorithm 2 Unconditional Hybrid Sampler for TimeBridge
Hyperparameter Setups
For experimental setups, we follow the base setups of Yuan and Qiao (2024). Therefore, we try to use the minimal search space for determining only for diffusion scheduler ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are the functions of times.
As demonstrated in EDM (Karras et al., 2022), we can reparameterize $\alpha_{t}$ and $\sigma_{t}$ using $\beta_{min}$ and $\beta_{d}$ as follows:
(31) $\alpha_{t}$ $=e^{-\frac{1}{4}\beta_{d}t^{2}-\frac{1}{2}\beta_{min}t}$
(32) $\sigma_{t}$ $=e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{min}t}-1,$
With the aforementioned equations, we can adjust $\alpha_{t}$ and $\sigma_{t}$ based on $\beta_{min}$ and $\beta_{d}$.
Furthermore, we can adjust the weight scheduler of $w_{t}$ in Equation (18) with $\sigma_{data}$ as in EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) as follows:
(33) $w_{t}=\frac{(\sigma_{t}^{2}+\sigma_{data}^{2})}{(\sigma_{t}\cdot\sigma_{data})^{2}}.$
Note that EDM (Karras et al., 2022) and DDBM (Zhou et al., 2024) uses this $\sigma_{data}$ in the reparmetriziation of $F$-matching in Equation (28) to calculate, $c_{skip}$, $c_{out}$, $c_{in}$, $c_{noise}$, but we exclude them since we use direct $D$-matching.
In the GP setups for the unconditional generation, we additionally search the kernel parameter $\eta$.
The summarized hyperparameters used for unconditional generation and conditional generation are demonstrated in Tables 11 and 12 for unconditional and conditional settings, respectively.
@@ B.4. Additional Experimental Details
Base setup (Yuan and Qiao, 2024) Batch size - 128 128 128 64 64 64
@@
Table 12. Hyperparameters for conditional generation. $\dagger$ We use 0.025 for the Energy dataset with Butterworth trend setting.
@@
We unified the sampling settings with the second-order Heun sampler with 40 steps, which have 119 NFE. We use the churn ratio $s=0.33$ for stochastic sampling in Algorithms 1 and 2.

### v3

@@ 7. Discussion and Conclusion
Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) (No. RS-2024-00338859), and the Institute of Information & communications Technology Planning & Evaluation (IITP) (No. RS-2022-II220984) grants funded by the Korea government (MSIT). Jinseong Park is supported by a KIAS Individual Grant (AP102301) via the Center for AI and Natural Sciences at Korea Institute for Advanced Study. This work was supported by the Center for Advanced Computation at Korea Institute for Advanced Study.
@@ A.1. Schrödinger Bridge
Standard diffusion models suffer from limitations such as the requirement of a sufficiently long time for the prior distribution to approximate a standard Gaussian distribution. To address these challenges, diffusion Schrödinger bridge models have been introduced (de2021diffusion; chen2021likelihood) to find an entropic optimal transport between two probability distributions in terms of Kullback–Leibler divergence on path spaces (chen2021likelihood; leonard2012schrodinger). While standard diffusion models and flow matching models are not guaranteed to provide optimal transport, Schrödinger bridge models aim to find paths that recover entropy-regularized versions of optimal transport (shi2024diffusion) as follows:
(21) $\small\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$
@@
where $\mathcal{P}_{[0,T]}$ refers to the space of path measures on $[0,T]$, $p^{ref}$ denotes the reference path measure, and $p_{0},p_{T}$ represents the marginal distributions of $p$ at each time step $0,T$, respectively. Setting the reference path as the forward SDE, the problem becomes equivalent to following forward-backward SDEs:
(22) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\end{split}$
(23) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are described as:
@@
(24) $\begin{split}&\frac{\partial\bm{\Psi}}{\partial t}=-\nabla_{x}\bm{\Psi}^{T}{\mathbf{f}}-\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\bm{\Psi})\\ &\frac{\partial\hat{\bm{\Psi}}}{\partial t}=-\nabla_{x}\cdot(\hat{\bm{\Psi}}{\mathbf{f}})+\frac{1}{2}tr(g^{2}\nabla^{2}_{x}\hat{\bm{\Psi}})\\ s.t.\quad\bm{\Psi}_{0}\hat{\bm{\Psi}}_{0}=&p_{data},\bm{\Psi}_{T}\hat{\bm{\Psi}}_{T}=p_{prior},p_{t}=\bm{\Psi}_{t}\hat{\bm{\Psi}}_{t}\ \text{for}\ t\in[0,T].\end{split}$
@@
The Schrödinger bridge problem can be addressed using iterative methods (kullback1968probability). At the outset, several early studies employed iterative methods to simulate trajectories converging to the Schrödinger bridge problem (vargas2021solving; shi2024diffusion; de2021diffusion). A classical solver for Schrödinger bridges is Iterative Proportional Fitting (IPF). Beginning with arbitrary half-potentials $(\bm{\Psi}^{(0)}_{t},\hat{\bm{\Psi}}^{(0)}_{t})$, the algorithm alternates two normalization steps:
(25) $\begin{split}\text{(Forward-update)}\quad\hat{\bm{\Psi}}^{(n+1)}_{t}=\frac{p_{data}}{\int\bm{\Psi}^{(n)}_{t}\,d{\mathbf{x}}_{0}},\\ \text{(Backward-update)}\quad\bm{\Psi}^{(n+1)}_{t}=\frac{p_{prior}}{\int\hat{\bm{\Psi}}^{(n+1)}_{t}\,d{\mathbf{x}}_{T}}.\end{split}$
@@
After each full iteration, the product $\bm{\Psi}^{(n+1)}_{t}\hat{\bm{\Psi}}^{(n+1)}_{t}$ matches both marginals, the sequence converges geometrically, requiring one forward–backward integration per iteration. However, these methods rely on expensive iterative approximation techniques and have seen limited empirical application (zhou2024denoising).
@@ A.2. Denoising Diffusion Bridge Models
Alternatively, chen2021likelihood leveraged the theory of forward-backward SDEs to derive an exact log-likelihood expression for the Schrödinger bridge, which accurately generalizes the approach for score generative models. Drawing inspiration from the perspective that the framework of diffusion bridges, employing forward-backward SDEs, encompasses the paradigms of score matching diffusion models (song2020denoising) and flow matching optimal transport paths (lipman2022flow). zhou2024denoising demonstrated that reintroducing several design choices from these domains becomes feasible. In particular, the reparameterization outlined in karras2022elucidating are utilized.
Table 13. VP and VE of diffusion bridge settings in DDBM.
@@
For noise scheduling, common options include variance-preserving (VP) and variance-exploding (VE) diffusion (song2020denoising). We express the reparameterization for the VP and VE bridge of zhou2024denoising in Table 13. In our study, we chose to utilize the VP bridge, as our experimental findings indicate its superiority of VP over VE. For implementing the diffusion sampler, we modify the official code of DDBM (zhou2024denoising) in https://github.com/alexzhou907/DDBM.
@@ Appendix B Experimental Settings
Base setup (yuan2024diffusionts) Batch size - 128 128 128 64 64 64
@@
B.1. Baseline methods
For unconditional generation, we compare with TimeVAE (desai2021timevae), TimeGAN (yoon2019time), Cot-GAN (xu2020cot), Diffwave (kong2021diffwave), DiffTime (coletta2024constrained), and Diffusion-TS (yuan2024diffusionts). For Table 2, we note the quality evaluation results in Diffusion-TS. We reimplement Diffusion-TS using the official GitHub in https://github.com/Y-debug-sys/Diffusion-TS, unfortunately, we cannot achieve the performance in the original paper. For conditional generation, we reimplement SSSD (lopez2023diffusionbased) using the official GitHub https://github.com/SSSD/sssd. For the trend-guided synthesis, we treat the trend as conditional observations and train the SSSD model to match the total time series data. We add the conditional embedding layers to their original model.
B.2. Evaluation Metrics
FID score is a widely used metric in image tasks, capturing the similarity between generated images and real ones (heusel2017gans). To adapt it to the time series domain, jeha2022psa proposed Context-FID score, which replaces the Inception model used in the original FID with TS2Vec, a pre-trained time series representation learning model (yue2022ts2vec). As Context-FID score aligns with the predictive accuracy of the generative model in downstream tasks, a lower Context-FID score indicates higher synthetic data quality. Both synthetic and real time-series data are encoded using a pre-trained TS2Vec model, and the FID score is calculated on the representations.
Correlational score is a metric used to evaluate the similarity in the covariance of time series features between original data and synthetic data (liao2020conditional). First, we estimate the covariance between the $i$-th and $j$-th features of the time series by:
Then compute the correlation between two datasets as follows:
B.3. Additional Experimental Details
Datasets
Table 15. Details of datasets.
Dataset Sample size # of features Datatset link
Sines 10000 5 https://github.com/jsyoon0823/TimeGAN
Stocks 3773 6 https://finance.yahoo.com/quote/GOOG
ETTh 17420 7 https://github.com/zhouhaoyi/ETDataset
MuJoCo 10000 14 https://github.com/google-deepmind/dm_control
Energy 19711 28 https://archive.ics.uci.edu/dataset
fMRI 10000 50 https://www.fmrib.ox.ac.uk/datasets
Table 16. Hyperparameters for conditional generation. $\dagger$ We use 0.025 for the Energy with the Butterworth trend.